An X ray tube operates at an applied voltage of  15000 v . what is the shortest wavelength of the X ray emitted by this tube?

Grade:

1 Answers

Arun
5 years ago

Where is Planck’s constant, is speed of light and eV the applied potential (for this purpose, the same as KV).

So

lambda(min) = hc/E = 6.6* 10^-34 * 3* 10^8 / 1.6 * 10^-19 * 15000

 = 0.825 * 10^-10 m